# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v2

@@ II.2 Dissociation Curves
We pursue with the dissociation energies reported in Figure 3. Clearly, the simple ADAPT-VQE/STO-3G level of theory appears quite far from an accurate description of the dissociation (compared to the calculations in the largest basis sets), which is achieved with FCI/triple-zeta (and beyond) levels. The non-corrected ADAPT-VQE values are represented with bold lines whereas the corrected values are reported with dashed curves. These dissociation curves are extremely challenging as they involve several regimes of correlation going from weak correlation happening typically at short distances to strong correlation effects at large distances. First, we notice that for all the cases, the basis-set corrected values around the equilibrium are always substantially closer to the large cc-pV5Z reference values than the uncorrected ones and improve with the basis-set size and the use of SABS. For larger distances, the basis-set requirements appear even more stringent. In practice, the three dissociation curves manage to be converged to high accuracy thanks to the use of SABS: we obtain nearly a triple-zeta quality for N2 using the basis-set correction with the V5Z-6 basis set, nearly a cc-pV5Z quality for H2 using the basis-set correction with the V5Z-8 basis set, and nearly a cc-pV5Z quality for LiH using the basis-set correction with a V5Z-7 basis set. For H2, the basis-set corrected cc-pVDZ curve matches the cc-pV5Z reference up to a distance of 2.5 Å and a slight discrepancy appears at long distances. This is consistent with classical computations leading to a convergence to the CBS limit only when the basis-set correction is applied to the cc-pVTZ basis set (114 qubits). It is possible to fix this issue more affordably by using a V5Z-8 basis set requiring only 24 qubits. For LiH, it is also possible to fix the small discrepancies observed on the V5Z-7 basis-set dissociation curve using a larger V5Z-10 basis set (28 qubits) which reaches the CBS limit.
@@
Finally, it is important to highlight the good performance of the DBBSC method for the prediction of the dissociation curve of the triple-bonded N2 molecule. Indeed, such computation is well-documented in the literature and known as extremely difficult as it requires both a multi-reference treatment and various weak correlation effects going from short-range to charge polarization effects [46, 47, 48, 49]. At the cost of a minimal basis set (i.e. 16 qubits), our N2 DBBSC computations achieved nearly a cc-pVTZ quality, providing an accuracy which would have required around 100 logical qubits in the context of a brute-force simulation. To the best of our knowledge, these results are the most accurate using a quantum algorithm when compared to the recent results from the literature. Indeed, a group of researchers managed to predict this dissociation curve using a Local Unitary Cluster Jastrow (LUCJ) ansatz coupled to double-zeta basis sets (6-31G and cc-pVDZ) [50]. The computation required the use of massive computational resources, namely hundreds of compute nodes of the Fugaku classical supercomputer coupled to a QPU. Alternatively, in the context of the fermionic quantum emulator, N2 simulations used ADAPT-VQE coupled to the 6-31G and def2-SVP (56 qubits) basis sets to perform computation via approximate Matrix Product States (MPS) [51]. In the present work, we achieve a better accuracy with far less qubits and, again, these results can be systematically improved using larger SABS. Indeed, going beyond a cc-pVTZ basis set is possible using the next (larger) SABS in term of size, i.e. V5Z-11 (32 qubits).
@@ III Conclusion
An additional reduction of the required basis size is provided by our SABS approach. Such a “black-box” pivoted-Cholesky strategy for the on-the fly generation of basis sets has been shown to be competitive and often superior to available standard choices. It offers systematic computational savings for large basis sets, reducing significantly the qubit requirements. Thus the SABS’ usefulness is not limited to QC algorithms as they also offer systematically improvable solutions for performing reference classical computations. Indeed, since SABS are user-defined truncated versions of the Dunning basis sets that can match a qubit budget, they also offer a reduced cost access to improved accuracy for any quantum-chemistry method.
@@
In this context, beside the state-of-the-art ADAPT-VQE hybrid quantum-classical algorithm, it would be interesting to revisit accurate fixed wave-function ansätze [56], such as UCCSDT [57] and others, to analyze their convergence when coupled to the DBBSC method. In practice, the presented basis-set correction framework is not restricted to a given QC wave-function ansatz and can leverage any future QC algorithmic improvements. Our present DBBSC/SABS methodology still requires to use qubits either through quantum hardware or through the use of a classical quantum emulator. The qubit count is therefore our main limitation. The computations of this paper used up to 32 qubits and represent a proof-of-concept study of what one can do with quantum emulation. However, state-vector simulations have limitations due to memory. Indeed, if they are theoretically possible up to 40-50 qubits on very large exascale supercomputers, they start to become relatively unpractical for quantum-chemistry simulations when reaching 36 qubits due to the high computational resources and time-to-solution requirements. To explore further the chemical electronic space with quantum algorithms, we are currently upgrading our Hyperion-1 framework to increase our emulated qubit counts by going beyond the state-vector formalism thanks to various elements from the CUDA-Q SDK [58, 59] developed by NVIDIA. We are also presently testing various implementations of the DBBSC algorithms on available quantum hardware. To conclude, by reducing the number of qubits required to reach the CBS limit, we expect to tackle predictive real-world quantum-chemistry applications with strategies applicable to both NISQ and FTQC algorithms.
@@ Acknowledgments
We thank the PEPR EPIQ - Quantum Software (ANR-22-PETQ-0007, J.-P.P.) for funding D. Traore’s postdoctoral position. I.-M. Lygatsika’s PhD position has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant No 810367, project EMC2, J.-P. P. and Y.M.). Computations have been performed at IDRIS (Jean Zay) on GENCI Grant no A0150712052 (J.-P.P.), at Scaleway (Jero) and on the SuperPOD reference cluster EOS at NVIDIA. The authors thank D. Ruau and C. Hundt (NVIDIA) for continuous support, as well as the NVIDIA Quantum team for technical discussions.

### v3

@@ II.2 Dissociation Curves
We pursue with the dissociation energies reported in Figure 3. Clearly, the simple ADAPT-VQE/STO-3G level of theory appears quite far from an accurate description of the dissociation (compared to the calculations in the largest basis sets), which is achieved with FCI/triple-zeta (and beyond) levels. The non-corrected ADAPT-VQE values are represented with bold lines whereas the corrected values are reported with dashed curves. These dissociation curves are extremely challenging as they involve several regimes of correlation going from weak correlation happening typically at short distances to strong correlation effects at large distances. First, we notice that for all the cases, the basis-set corrected values around the equilibrium are always substantially closer to the large cc-pV5Z reference values than the uncorrected ones and improve with the basis-set size and the use of SABS. For larger distances, the basis-set requirements appear even more stringent. In practice, the three dissociation curves manage to be converged to high accuracy thanks to the use of SABS: we obtain nearly a triple-zeta quality for N2 using the basis-set correction with the V5Z-6 basis set(16 qubits). A triple-zeta-like quality can be achieved using the larger V5Z-11 basis-set at the price of more qubits (32 qubits, see Table II). In the same line, nearly a cc-pV5Z quality for H2 using the basis-set correction with the V5Z-8 basis set, and nearly a cc-pV5Z quality for LiH using the basis-set correction with a V5Z-7 basis set. For H2, the basis-set corrected cc-pVDZ curve matches the cc-pV5Z reference up to a distance of 2.5 Å and a slight discrepancy appears at long distances. This is consistent with classical computations leading to a convergence to the CBS limit only when the basis-set correction is applied to the cc-pVTZ basis set (114 qubits). It is possible to fix this issue more affordably by using a V5Z-8 basis set requiring only 24 qubits. For LiH, it is also possible to fix the small discrepancies observed on the V5Z-7 basis-set dissociation curve using a larger V5Z-10 basis set (28 qubits) which reaches the CBS limit.
@@
Finally, it is important to highlight the good performance of the DBBSC method for the prediction of the dissociation curve of the triple-bonded N2 molecule. Indeed, such computation is well-documented in the literature and known as extremely difficult as it requires both a multi-reference treatment and various weak correlation effects going from short-range to charge polarization effects [80, 81, 82, 83]. At the cost of a minimal basis set (i.e. 16 qubits), our N2 DBBSC computations achieved nearly a cc-pVTZ quality, providing an accuracy which would have required around 100 logical qubits in the context of a brute-force simulation. To the best of our knowledge, these results are the most accurate using a quantum algorithm when compared to the recent results from the literature. Indeed, a group of researchers managed to predict this dissociation curve using a Local Unitary Cluster Jastrow (LUCJ) ansatz coupled to double-zeta basis sets (6-31G and cc-pVDZ) [84]. The computation required the use of massive computational resources, namely hundreds of compute nodes of the Fugaku classical supercomputer coupled to a QPU. Alternatively, in the context of the fermionic quantum emulator, N2 simulations used ADAPT-VQE coupled to the 6-31G and def2-SVP (56 qubits) basis sets to perform computation via approximate Matrix Product States (MPS) [85]. In the present work, we achieve a better accuracy with far less qubits and, again, these results can be systematically improved using larger SABS. Indeed, access to a cc-pVTZ-like basis set accuracy is possible using the next (larger) SABS in term of size, i.e. V5Z-11 (32 qubits).
@@ III Conclusion
An additional reduction of the required basis-set size is provided by our SABS approach. Beside being fast, as SABS can be generated within seconds, such a “black-box” pivoted-Cholesky strategy for the on-the fly generation of basis sets has been shown to be competitive and often superior to available standard choices. It offers systematic computational savings for large basis sets, reducing significantly the qubit requirements. Thus the SABS’ usefulness is not limited to QC algorithms as they also offer systematically improvable solutions for performing reference classical computations. Indeed, since SABS are user-defined truncated versions of the Dunning basis sets that can match a qubit budget, they also offer a reduced cost access to improved accuracy for any quantum-chemistry method.
@@
In this context, beside the state-of-the-art ADAPT-VQE hybrid quantum-classical algorithm, it would be interesting to revisit accurate fixed wave-function ansätze [89], such as UCCSDT [90] and others, to analyze their convergence when coupled to the DBBSC method. In practice, the presented basis-set correction framework is not restricted to a given QC wave-function ansatz and can leverage any future QC algorithmic improvements. Our present DBBSC/SABS methodology still requires to use qubits either through quantum hardware or through the use of a classical quantum emulator. The qubit count is therefore our main limitation. The computations of this paper used up to 32 logical qubits and represent a proof-of-concept study of what one can presently do with quantum emulation to prepare the advent of fault-tolerant quantum computing [91], see Table IV (Appendix) for detailed resource estimations. However, state-vector simulations have limitations due to memory. Indeed, if they are theoretically possible up to 40-50 qubits on very large exascale supercomputers, they start to become relatively unpractical for quantum-chemistry simulations when reaching 36 qubits due to the high computational resources and time-to-solution requirements. To explore further the chemical electronic space with quantum algorithms, we are currently upgrading our Hyperion-1 framework to increase our emulated qubit counts by going beyond the state-vector formalism thanks to various elements from the CUDA-Q SDK [92, 93] developed by NVIDIA. We are also presently testing various implementations of the DBBSC algorithms on available quantum hardware. To conclude, by reducing the number of qubits required to reach the CBS limit, we expect to tackle predictive real-world quantum-chemistry applications with strategies applicable to both NISQ and FTQC algorithms.
@@ Acknowledgments
We thank the PEPR EPIQ - Quantum Software (ANR-22-PETQ-0007, J.-P.P.) for funding D. Traore’s initial postdoctoral position. Support of the HQI initiative is also acknowledged (J.-P. P. and Y.M.). I.-M. Lygatsika’s PhD position has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant No 810367, project EMC2, J.-P. P. and Y.M.). Computations have been performed at IDRIS (Jean Zay) on GENCI Grants: no A0150712052 (J.-P. P.) and grant GC010815453 (Grand Challenge H100 Jean Zay, J.-P. P.), at Scaleway (Jero) and on the SuperPOD reference cluster EOS at NVIDIA. The authors thank D. Ruau and C. Hundt (NVIDIA) for continuous support, as well as the NVIDIA Quantum team for technical discussions.
Appendix: CNOT counts
Table 4: CNOT-gate counts for qubit and qubit-excitation-based (QEB) operator pools. The numbers of CNOT gates, $N_{\text{CNOT}}$, are evaluated as $N_{\text{single}}^{3}+13N_{\text{double}}$ for the QEB pool, and as $N_{\text{single}}^{2}+6N_{\text{double}}$ for the qubit pool, where $N_{\text{simple}}$ is the number of single-qubit operators and $N_{\text{double}}$ is the number of two-qubit operators. $N_{\text{op}}$ is the number of operators in the final wave-function ansatz, and $N_{\text{adapt iter}}$ is the number of ADAPT-VQE iterations achieved and at which we collect the values in the table.
H2 $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 4 13 6 1 1
6-31G 8 71 34 7 7
cc-pVDZ 20 233 110 21 21
V5Z-8 24 395 186 35 35
H4 $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 8 207 98 19 19
H6 $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 12 2437 1134 199 199
H8 $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 16 12677 5870 999 999
6-31G 32 8855 4090 685 685
He $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
pc-seg0 4 19 10 3 3
6-31G 4 19 10 3 3
cc-pVDZ 10 58 28 6 6
Be $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 8 39 18 3 3
pc-seg0 10 58 28 6 6
6-31G 16 207 98 19 19
cc-pVDZ 26 292 136 26 26
LiH $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 10 90 44 10 10
pc-seg0 14 258 124 26 26
6-31G 20 511 242 47 47
VQZ-4 10 90 44 10 10
V5Z-4 10 90 44 10 10
V5Z-7 16 291 138 27 27
V5Z-10 28 1083 506 91 91
H2O $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 12 729 342 63 63
6-31G 24 12657 5862 999 999
V5Z-11 30 12647 5858 999 999
N2 $N_{\text{qubits}}$ $N_{\text{CNOT}}$ (if QEB) $N_{\text{CNOT}}$ (if qubit) $N_{\text{op}}$ $N_{\text{adapt iter}}$
STO-3G 16 4868 2256 386 386
V5Z-6 16 9714 4492 758 758
V5Z-11 32 11718 5420 916 916